# QG-SMS: Enhancing Test Item Analysis via Student Modeling and Simulation

### v1

@@ 1 Introduction
Figure 1: Existing LLM-based approaches rely solely on question content for evaluation. In this example, ChatEval identifies $Q_{1}$ as the better test item for distinguishing high- and low-performing students, reasoning that it requires learners to apply a concept rather than merely recall information (as in $Q_{2}$). However, real student performance data shows $Q_{1}$ has lower discrimination. This highlights the need for evaluation methods that incorporate student modeling.
In this work, we introduce test item analysis, a well-established method in education for assessing test item quality, into the QG evaluation pipeline. In educational testing, test item quality is assessed through both pre-examination and post-examination analyses. Pre-examination analysis evaluates test items (i.e., quiz questions) before administration, focusing on dimensions such as topic alignment, where instructors or subject matter experts ensure that test content aligns with learning objectives Mahjabeen et al. (2017). Post-examination analysis is a powerful tool that evaluates the quality of test questions by analyzing how test takers respond to them. It occurs after test administration, providing insights into dimensions such as item difficulty, item discrimination, and distractor efficiency through statistical analyses of test-taker performance Mahjabeen et al. (2017). Post-examination analysis can help improve the test’s validity and reliability, which is valuable for improving test items that will be used again in later tests. However, it cannot evaluate test questions during the test design phase, as it requires test-taker responses which are only available after the test has been administered.
@@
Figure 2: Performance of LLM-based evaluation methods (defined in §$\S$4.2) in pairwise test item comparisons. Existing approaches (in colors except purple/markers except stars) perform well in pre-examination analysis (95.6% on average). However, their post-examination performance on question difficulty, discrimination, and distractor efficiency, significantly falls behind, with average consistent accuracies of 49.1%, 44.5%, and 53.3%, respectively. Our proposed approach, QG-SMS, bridges this gap, outperforming all methods across all dimensions.
Fig. 1 illustrates the shortcomings of existing LLM-based evaluation approaches for post-examination analysis. These methods primarily assess question content while neglecting test-taker perspectives, which are crucial for evaluating question quality. To address this gap, we propose QG-SMS, a novel evaluation framework (illustrated in Fig.  3) that utilizes a large language model (LLM) to simulate students with diverse levels of understanding for test item analysis. These simulations serve as reliable indicators of student performance on candidate test items, significantly enhancing the LLM’s capacity for evaluating question quality (Fig.  2). In summary, this paper makes the following contributions:
@@ 2 Problem Definition
Figure 3: QG-SMS follows three steps: (1) Generating student profiles with diverse understanding of learning materials, (2) Predicting responses of simulated students to candidate questions, and (3) Evaluating question quality based on simulated student performance. In the same example shown in Fig. 1, QG-SMS arrives at the opposite conclusion from existing evaluation approaches. According to the simulation, applications of computer vision (covered in $Q_{1}$) are common knowledge among students, including Alice - The Attentive and Bob - The Beginner, making them equally likely to provide a correct response. Meanwhile, recalling a specific statistic from the lecture (as required by $Q_{2}$) targets students who pay closer attention like Alice - The Attentive. Based on the simulated performance, QG-SMS correctly identifies $Q_{2}$ as the question with higher discrimination.
@@ 2.2 Task Definition
Given learning materials $L$ such as lecture content or transcripts, our goal is to obtain a test question that effectively assesses students’ knowledge of $L$. Since instructors may have varying requirements for test questions Wang et al. (2022a), let $R_{d}$ denote the desired characteristic or requirement of a test question with respect to a specific dimension $d$ such as question difficulty, discrimination, topic coverage, or distractor efficiency. Given two candidate questions $Q_{1}$ and $Q_{2}$ derived from $L$, the task is to determine which question better satisfies the requirement $R_{d}$ [^1] While the current task setup relies on binary comparisons, an extended approach using multiple pairwise comparisons could establish a ranking-based system, where question rankings translate into computed DF/DE/DC scores.. We provide an example of the task in Fig. 1.
@@ 3 QG-SMS: Student Modeling and Simulation for Test Item Analysis
During the test design phase, it is imperative to anticipate the diverse ways students may interpret learning materials. For example, in multiple-choice tests, effective distractors help teachers identify students who hold certain misconceptions Gierl et al. (2017). In this sense, to enhance the educational alignment of automated test item evaluation, we propose QG-SMS, which leverages LLM to model and simulate how well test items measure varying levels of student understanding. As illustrated in Fig. 3, QG-SMS consists of three key steps: (1) student profile generation, (2) student performance prediction, and (3) evaluation.
@@
Step 3 - Evaluation: Finally, QG-SMS assesses whether a test item fulfills its intended purpose by examining the responses of students with different levels of understanding. For example, an easy question should yield correct answers from a wide range of students, while a challenging question should only be correctly answered by those who have a deeper understanding of the learning materials. Formally, given the pair of candidate questions $\{Q_{1},Q_{2}\}$, the desired characteristic of the test item $R_{d}$ and the predicted student performance from step 2, the task is to determine which question better satisfies requirement $R_{d}$.
@@ 4 Experiments
Method Difficulty (DF) Discrimination (DC) Dist. Eff. (DE)
EduAgent 124 pairs DBE-KT 162 pairs EduAgent 61 pairs DBE-KT 93 pairs EduAgent 75 pairs
AA CA AA CA AA CA AA CA AA CA
Individual Scoring
BERTScore 51.61 - 61.73 - 65.57 - 30.11 - 65.33 -
KDAlarge 60.48 - 54.32 - 60.66 - 58.06 - 77.33 -
Pairwise LLM-based
Vanilla 63.71 50.80 67.28 49.38 63.11 49.18 63.98 49.46 73.33 64.00
CoT 61.69 32.26 64.20 38.89 59.84 32.79 62.90 34.41 60.00 28.00
Metrics 65.32 53.22 64.20 48.77 65.57 50.82 61.29 45.16 72.00 62.67
Reference 66.53 51.61 62.96 45.06 62.30 45.90 60.75 44.09 69.33 56.00
Swap 66.53 54.84 68.31 53.70 64.75 45.90 62.90 48.39 68.00 53.33
ChatEval 68.95 51.61 70.99 59.88 54.92 42.56 65.05 53.76 69.33 56.00
QG-SMS (Ours) 68.55 65.32 69.44 64.20 66.39 55.74 66.66 56.99 79.33 74.67
Table 1: Performance (AA: average accuracy, CA: consistent accuracy) of existing QG evaluation approaches and our proposed QG-SMS approach in test item analysis, grouped by dimension and dataset. The highest and second-highest values for each column are highlighted with bold and underline markers, respectively.
@@ 4.2 QG Evaluators
We compare QG-SMS with two individual-scoring metrics: the reference-based BERTScore Zhang et al. (2019) and the reference-free KDA Moon et al. (2022). For BERTScore, since we do not have a reference question for each pair, we instead use the learning material $L$ as the reference and measure the similarity between $L$ and each question. For KDA, we use the large version of the model-based metric. As these metrics assign separate scores to $Q_{1}$ and $Q_{2}$, we must determine how to compare their scores to establish a preference. For each dimension, we select the direction that yields the highest average accuracy for the EduAgent dataset:
  • Easier question: $\downarrow$ BERTScore, $\uparrow$ KDA
  • Higher discrimination: $\uparrow$ BERTScore, $\downarrow$ KDA
  • Higher distractor efficiency:$\uparrow$ BERTScore, $\downarrow$ KDA
We retain this comparison direction for the DBE-KT dataset, as a reliable metric should exhibit consistent behavior across domains.
@@ 5.1 Enhancing Test Item Analysis with QG-SMS
Reference-based metrics like BERTScore are not reliable in reflecting the educational value of test items, as their evaluation behavior for the same dimension varies significantly across domains. Tbl. 1 highlights this inconsistency: when selecting the question with the higher BERTScore as the question with higher discrimination, the average accuracy for the EduAgent dataset (Introduction to AI lectures) is 65.57%. Meanwhile, for the DBE-KT dataset (Relational Database exercises), the accuracy within the same domain drops to 30.11%.
Beyond reference-based metrics, existing LLM-based QG evaluation approaches also struggle with post-examination analysis, as shown in Fig. 2. To address this gap, QG-SMS enhances test item analysis performance by incorporating student modeling and simulation, as demonstrated in Tbl. 1. Across both datasets, QG-SMS achieves the highest average accuracy in evaluating DC and DE, and the second-highest average accuracy in evaluating DF. Additionally, QG-SMS significantly outperforms all baselines in consistent accuracy, demonstrating its robustness to input order variations. For instance, QG-SMS’s consistent accuracy for DF in the EduAgent dataset is $65.32\%$, maintaining a $10.48\%$ gap over the second-best baseline (Swap). Fig. 3 provides a case study illustrating how simulation enhances test item analysis, facilitating a more educationally aligned evaluation.
@@ 5.2 Analysis
Varying $\alpha$: We further investigate the effectiveness of QG-SMS compared to other LLM-based approaches across different values of $\alpha$, i.e., the threshold of quality difference between a pair of questions. Fig. 4 indicates that the performances of all LLM-based metrics consistently improve as $\alpha$ increases. This trend is intuitive, as higher $\alpha$ values suggest a larger quality gap between question pairs, making the evaluation task easier. Importantly, QG-SMS remains the top performer regardless of the changes in $\alpha$.
@@ 5.3 Human Evaluation Study
Study Description: We recruit three volunteer annotators, including two graduate and one undergraduate student in Computer Science. Their domain knowledge is highly related to the lecture contents of the EduAgent dataset (e.g., AI related knowledge) and they all have some teaching experience. Annotators are tasked to make preferences on $120$ pairs of questions, including $60$ pairs of human-written and $60$ pairs of machine-generated questions. Each pair differs in one of three dimensions - DF, DC, and DE. We use the EduAgent dataset. Its lectures target a general audience, supporting the credibility of our annotators in assessing lecture content and quiz questions. We provide more details on the question generation process and instructions given to annotators in §$\S$§A.3.
Study Results: In 75 of 120 cases (62.5%) all three annotators agree on the same preference. For the remaining cases, we adopt the majority preference (chosen by 2 out of 3 annotators) as the representative of human judgment. We report the results of our human evaluation study in Tbl. 2.
In human-written question pairs with ground-truth labels based on student performance, our human annotators achieve the highest average accuracy ($78.33\%$) compared to LLM-based evaluators. When broken down by dimension, the average accuracy of human annotators is $90.48\%$, $53.33\%$, and $87.5\%$ for DF, DC, and DE respectively. This observation suggests that performing item analysis on the DC dimension poses significant challenges to our annotators. As they noted during post-examination feedback, it is challenging to identify which question more effectively distinguishes between high-performing and low-performing students when they do not have access to the specific student profiles in the classroom. In terms of evaluating DC, our proposed QG-SMS surpasses human annotators, and on the other two dimensions, DF and DE, QG-SMS achieves the closest accuracy scores to humans. On average, QG-SMS achieves the second-highest accuracy—surpassed only by human annotators. The results show the effectiveness of simulating student understanding and performance. See Tbl. 3 for detailed results.
For the other 60 pairs of generated questions, we use the human annotators’ preferences as the labels and evaluate the performance of QG evaluators accordingly. It can be seen from Tbl. 2 that QG-SMS achieves the highest average accuracy and consistent accuracy in this setting, demonstrating state-of-the-art alignment with human judgment.

### v2

@@ 1 Introduction
Learning Material
Introduction of computer vision: Computer vision (CV) is the field of computer science that focuses on creating digital systems that can process, analyze, and make sense of visual data […]. For example, […]
Computer vision history […] In 2012, a team from the University of Toronto […]. The model, called AlexNet, […],achieved an error rate of 16.4%, which overperformed all other methods at that time. […]
Quiz Questions
$Q_{1}$: Which of the following may utilize computer vision techniques? 1). Use a camera to check potential issues on the surface of products (2). Estimate the freshness of apples from pictures (3). Estimate whether a car is speeding via a camera (4). Determine whether a piece of audio is spoken by a specific person
A) (1)(2)(3); B) (1)(2)(4); C) (2)(3)(4); D) (1)(2)(3)(4).
$Q_{2}$: One breakthrough in computer vision happened at the University of Toronto in 2012, which achieved an error rate of [ ] in image classification.
A) 6.4%; B) 10.4% C) 12.4% D) 16.4%.
Evaluation Task: Which question has higher discrimination?
Existing approaches: $Q_{1}$. $Q_{1}$ is an apply-level question, while $Q_{2}$ is a recall-level question.
Label based on Actual Student Performance: $Q_{2}$. Applications of CV appearing in $Q_{1}$ can be considered common knowledge while $Q_{2}$ tests a specific detail which only students who pay close attention to details may be able to answer.
Table 1: Existing LLM-based approaches rely solely on question content for evaluation. In this example, ChatEval identifies $Q_{1}$ as the better test item for distinguishing high- and low-performing students, reasoning that it requires learners to apply a concept rather than merely recall information (as in $Q_{2}$). However, real student performance data shows $Q_{1}$ has lower discrimination. This highlights the need for evaluation methods that incorporate student modeling. The complete case study is provided in Appendix A.4.
@@
In this work, we introduce test item analysis, a well-established method in education for assessing test item quality, into the QG evaluation pipeline. In educational testing, test item quality is assessed through both pre-examination and post-examination analyses. Pre-examination analysis evaluates test items (i.e., quiz questions) before administration, focusing on dimensions such as topic alignment, where instructors or subject matter experts ensure that test content aligns with learning objectives Mahjabeen et al. (2017). Post-examination analysis is a powerful tool that evaluates the quality of test questions by analyzing how test takers respond to them. It occurs after test administration, providing insights into dimensions such as item difficulty, item discrimination, and distractor efficiency through statistical analyses of test-taker performance Mahjabeen et al. (2017). Post-examination analysis can help improve future test items’ validity and reliability. However, it cannot evaluate test questions during the test design phase, as it requires test-taker responses that are only available after the test has been administered.
@@
Figure 1: Performance of LLM-based evaluation methods (defined in §$\S$4.2) in pairwise test item comparisons on the EduAgent dataset. Existing approaches (in colors except purple/markers except stars) perform well in pre-examination analysis (95.6% on average). However, their post-examination performance on question difficulty, discrimination, and distractor efficiency, significantly falls behind, with average consistent accuracies of 49.1%, 44.5%, and 53.3%, respectively. Our proposed approach, QG-SMS, bridges this gap, outperforming all methods across all dimensions.
Tbl. 1 illustrates the shortcomings of existing LLM-based evaluation approaches for post-examination analysis. These methods primarily assess question content while neglecting test-taker perspectives, which are crucial for evaluating question quality. To address this gap, we propose QG-SMS, a novel evaluation framework (illustrated in Fig.  2) that utilizes a large language model (LLM) to simulate students with diverse levels of understanding for test item analysis. These simulations serve as reliable indicators of student performance on candidate test items, significantly enhancing the LLM’s capacity for evaluating question quality (Fig.  1). In summary, this paper makes the following contributions:
@@
We release all implementation details of QG-SMS to facilitate future works  [^1] https://github.com/bnguyen5/qg-sms.
@@ 2 Problem Definition
Figure 2: QG-SMS follows three steps: (1) Generating student profiles with diverse understanding of learning materials, (2) Predicting responses of simulated students to candidate questions, and (3) Evaluating question quality based on simulated student performance. In the same example shown in Tbl. 1, QG-SMS arrives at the opposite conclusion from existing evaluation approaches. According to the simulation, applications of computer vision (covered in $Q_{1}$) are common knowledge among students, including Alice - The Attentive and Bob - The Beginner, making them equally likely to provide a correct response. Meanwhile, recalling a specific statistic from the lecture (as required by $Q_{2}$) targets students who pay closer attention like Alice - The Attentive. Based on the simulated performance, QG-SMS correctly identifies $Q_{2}$ as the question with higher discrimination.
@@ 2.2 Task Definition
Given learning materials $L$ such as lecture content or transcripts, our goal is to obtain a test question that effectively assesses students’ knowledge of $L$. Since instructors may have varying requirements for test questions Wang et al. (2022a), let $R_{d}$ denote the desired characteristic or requirement of a test question with respect to a specific dimension $d$ such as question difficulty, discrimination, topic coverage, or distractor efficiency. Given two candidate questions $Q_{1}$ and $Q_{2}$ derived from $L$, the task is to determine which question better satisfies the requirement $R_{d}$ [^2] While the current task setup relies on binary comparisons, an extended approach using multiple pairwise comparisons could establish a ranking-based system, where question rankings translate into computed DF/DE/DC scores.. We provide an example of the task in Tbl. 1.
@@ 3 QG-SMS: Student Modeling and Simulation for Test Item Analysis
During the test design phase, it is imperative to anticipate the diverse ways students may interpret learning materials. For example, in multiple-choice tests, effective distractors help teachers identify students who hold certain misconceptions Gierl et al. (2017). In this sense, to enhance the educational alignment of automated test item evaluation, we propose QG-SMS, which leverages LLM to model and simulate how well test items measure varying levels of student understanding. As illustrated in Fig. 2, QG-SMS consists of three key steps:
@@
Step 3 - Evaluation: Finally, QG-SMS assesses whether a test item fulfills its intended purpose by examining the responses of students with different levels of understanding. For example, an easy question should yield correct answers from a wide range of students, while a challenging question should only be correctly answered by those who have a deeper understanding of the learning materials. In this step, we leverage the LLM’s understanding of the question content along with the simulated performance data to make informed judgments on questions. Formally, given the pair of candidate questions $\{Q_{1},Q_{2}\}$, the desired characteristic of the test item $R_{d}$ and the predicted student performance from step 2, the task is to determine which question better satisfies requirement $R_{d}$.
@@ 4.2 QG Evaluators
We compare QG-SMS with three individual-scoring metrics:
The reference-based BERTScore Zhang et al. (2019) measures the semantic similarity between the candidate question and a reference. Since we do not have a reference question for each pair, we instead use the learning material $L$ as the reference and measure the similarity between $L$ and each question.
The reference-free KDA Moon et al. (2022) evaluates question quality based on the performance of simulated students with and without access to learning material $L$. We use the large version of this model-based metric.
The LLM-based QSalience Wu et al. (2024) measures the importance of the candidate question for understanding the learning material $L$. We use the best-performing model, mistral-instruct, as reported by its authors.
As these metrics assign separate scores to $Q_{1}$ and $Q_{2}$, we must determine how to compare their scores to establish a preference. For each dimension, we select the direction that yields the highest average accuracy for the EduAgent dataset (see Appendix A.2 for more details) and retain this comparison direction for the DBE-KT dataset, as a reliable metric should exhibit consistent behavior across domains.
Method Topic Coverage (TC) Difficulty (DF) Discrimination (DC) Dist. Eff. (DE)
EduAgent 217 pairs DBE-KT 286 pairs EduAgent 124 pairs DBE-KT 162 pairs EduAgent 61 pairs DBE-KT 93 pairs EduAgent 75 pairs
AA CA AA CA AA CA AA CA AA CA AA CA AA CA
Individual Scoring
BERTScore 79.26 - 40.20 - 51.61 - 61.73 - 65.57 - 30.11 - 65.33 -
KDAlarge 57.60 - 38.46 - 60.48 - 54.32 - 60.66 - 58.06 - 77.33 -
QSalience 54.84 - 48.25 - 54.03 - 60.49 - 52.46 - 47.31 - 68.00 -
Pairwise LLM-based
Vanilla 96.54 95.39 74.30 68.89 63.71 50.80 67.28 49.38 63.11 49.18 63.98 49.46 73.33 64.00
CoT 95.39 92.63 78.15 65.03 61.69 32.26 64.20 38.89 59.84 32.79 62.90 34.41 60.00 28.00
Metrics 97.70 97.70 80.59 75.17 65.32 53.22 64.20 48.77 65.57 50.82 61.29 45.16 72.00 62.67
Reference 97.00 96.31 72.55 66.43 66.53 51.61 62.96 45.06 62.30 45.90 60.75 44.09 69.33 56.00
Swap 95.85 95.85 81.64 74.48 66.53 54.84 68.31 53.70 64.75 45.90 62.90 48.39 68.00 53.33
ChatEval 96.77 95.85 80.94 74.13 68.95 51.61 70.99 59.88* 54.92 42.56 65.05 53.76 69.33 56.00
QG-SMS (Ours) 98.85 98.62 79.90 74.82 68.55 65.32* 69.44 64.20* 66.39 55.74 66.66 56.99 79.33 74.67*
Table 2: Performance (AA: average accuracy, CA: consistent accuracy) of existing QG evaluation approaches and our proposed QG-SMS approach in test item analysis, grouped by dimension and dataset. The highest and second-highest values for each column are highlighted with bold and underline markers, respectively. Asterisks (*) indicate statistical significance at p < 0.1 of LLM-based evaluation approach in improving CA against Vanilla.
@@ 5.1 Enhancing Test Item Analysis with QG-SMS
We provide insights into which dimensions of test item analysis that single-scoring metrics align the most closely, with the expectation that they should achieve accuracy $>50\%$ on both datasets. As shown in the Tbl. 2, their evaluation behavior is consistent for the DF dimension (i.e., easier questions tend to receive lower BERTScores, higher KDA values, and lower QSalience). KDA is also consistent in its evaluation for DC (higher discrimination questions tend to have lower KDA value), although the evaluation performance is not as comparable to QG-SMS (ours). However, the behavior of BERTScore and QSalience in TC and DC, and KDA in TC, appears dataset-specific and therefore not reliable in reflecting these educational aspects of test items.
While existing LLM-based evaluation approaches perform well in pre-examination analysis of topic coverage (TC), they struggle with post-examination dimensions, as shown in Fig. 1 and Tbl. 2. To address this gap, QG-SMS enhances test item analysis performance by incorporating student modeling and simulation. Across both datasets, QG-SMS achieves the highest average accuracy in evaluating DC and DE, and the second-highest average accuracy in evaluating DF. Additionally, QG-SMS significantly outperforms all baselines in consistent accuracy, demonstrating its robustness to input order variations. For instance, QG-SMS’s consistent accuracy for DF in the EduAgent dataset is $65.32\%$, maintaining a $10.48\%$ gap over the second-best baseline (Swap). Fig. 2 provides a case study illustrating how simulation enhances test item analysis, facilitating a more educationally aligned evaluation.
@@ 5.2 Analysis
Varying $\alpha$: We examine the effectiveness of QG-SMS compared to other LLM-based approaches across different values of $\alpha$, i.e., the threshold of quality difference in a pair of questions. Fig. 3 indicates that the performance of all LLM-based metrics consistently improves as $\alpha$ increases. This trend is intuitive, as higher $\alpha$ values suggest a larger quality gap in question pairs, making the evaluation task easier. Importantly, QG-SMS remains the top performer regardless of the changes in $\alpha$.
@@
Necessity of LLM-based evaluation step: In Step 3, we leverage the same LLM to make preference between candidate question pairs, providing the model with two inputs: the questions’ content, and the simulated student performance derived from Step 2. We believe that by providing the LLM with simulated student performances as augmented context, it can more effectively utilize its understanding of the questions’ semantics and nuanced language to make informed judgments. To assess the necessity of this step, we perform an ablation study in which we directly compute the Discrimination and Difficulty Index (DC and DF) using the simulated student performance and then compare the question pair based on these statistics. Results indicate that while this direct calculation yields better evaluation performance on Difficulty (with average accuracy improving from $68.55$ to $73.11$), it greatly harms Discrimination performance (with average accuracy decreasing from $66.39$ to $56.83$). This observation demonstrates that Step 3 is effective and necessary for the QG-SMS framework.
Extension of QG-SMS to questions without significant quality difference: We provide this analysis to demonstrate that our pipeline can be applied for test items with similar quality levels. Specifically, we apply QG-SMS to the Distractor Efficiency (DE) dimension within the EduAgent dataset, considering all possible question pairs with $\alpha$ of 00, $1$, $2$, or $3$. This setting results in a total of 308 pairs. For each unique question, we computed a ranking score as follows: If the question is consistently preferred in both swapped and non-swapped versions of a pair, we add 1 to its score; if it is preferred in only one version, we add 0.5. We then normalize the final score by the total number of pairs the question appears in.
Tbl. 3 compares QG-SMS-derived scores with actual DE values. An ANOVA test ($p<0.01$) reveals a significant difference in ranking scores among groups of questions with different DE levels. This supports our claim that QG-SMS can effectively identify groups of questions with similar quality on a specific dimension. We also compare the correlation between these ranking scores and actual DE values across methods. Tbl. 4 shows that QG-SMS achieves the highest correlation compared to the two strongest DE baselines (Vanilla and KDA).
$\text{DE}=0$ $\text{DE}=1$ $\text{DE}=2$ $\text{DE}=3$
0.22 $\pm$ 0.18 0.40 $\pm$ 0.22 0.48 $\pm$ 0.22 0.67 $\pm$ 0.22
Table 3: QG-SMS-derived ranking scores (mean and standard deviation) of questions with different distractor efficiency (DE) in the EduAgent dataset. Questions with higher DE tend to have higher QG-SMS ranking score.
Method Spearman Kendall Pearson
KDA 0.43 0.33 0.43
Vanilla 0.34 0.27 0.35
QG-SMS 0.48 0.38 0.50
Table 4: Correlation between method for computing ranking score and actual DE value of questions in the EduAgent dataset.
@@ 5.3 Human Evaluation Study
Study Description: We recruit three volunteer annotators, including two graduate and one undergraduate student in Computer Science. Their domain knowledge is highly related to the lecture contents of the EduAgent dataset (e.g., AI related knowledge) and they all have some teaching experience. Annotators are tasked to make preferences on $120$ pairs of questions, including $60$ pairs of human-written and $60$ pairs of machine-generated questions. Each pair differs in one of three dimensions - DF, DC, and DE. We use the EduAgent dataset. Its lectures target a general audience, supporting the credibility of our annotators in assessing lecture content and quiz questions. We provide more details on the question generation process and instructions given to annotators in Appendix A.3.
Study Results: In 75 of 120 cases (62.5%) all three annotators agree on the same preference. For the remaining cases, we adopt the majority preference (chosen by 2 out of 3 annotators) as the representative of human judgment. We report the results of our human evaluation study in Tbl. 5.
In human-written question pairs with ground-truth labels based on student performance, our human annotators achieve the highest average accuracy ($78.33\%$) compared to LLM-based evaluators. When broken down by dimension, the average accuracy of human annotators is $90.48\%$, $53.33\%$, and $87.5\%$ for DF, DC, and DE respectively. This observation suggests that performing item analysis on the DC dimension poses significant challenges to our annotators. As they noted during post-examination feedback, it is challenging to identify which question more effectively distinguishes between high-performing and low-performing students when they do not have access to the specific student profiles in the classroom. In terms of evaluating DC, our proposed QG-SMS surpasses human annotators, and on the other two dimensions, DF and DE, QG-SMS achieves the closest accuracy scores to humans. On average, QG-SMS achieves the second-highest accuracy—surpassed only by human annotators. The results show the effectiveness of simulating student understanding and performance. See Tbl. 7 for detailed results.
For the other 60 pairs of generated questions, we use the human annotators’ preferences as the labels and evaluate the performance of QG evaluators accordingly. It can be seen from Tbl. 5 that QG-SMS achieves the highest average accuracy and consistent accuracy in this setting, demonstrating state-of-the-art alignment with human judgment.